If one side of a 90 degree triangle is 32ft and the other is 18 ft how do you find the length of the 3rd side?
email address rms0617@aol.com

Answer by checkley77(12844)

If 32 & 18 are the 2 sides & you need the hypotenuse then:
32^2+18^2=c^2
1024+324=c^2
c^2=1348
c=sqrt1348
c=36.715 ans. for the hypotenuse.
--------------------------------------
If 32 is the hypotenuse then:
18^2+x^2=32^2
324+x^2=1024
x62=1024-324
x^2=700
x=sqrt700
x=26.4575 ans.